OpenStudy (anonymous):

Length of an arc in a circle of radius 12cm is 4 pi cm now find the size of the angle which describes the arc. explain me that plz

OpenStudy (anonymous):

 so, 4pi is the arclength, which is a fraction of the Circumference (proportional to the the angle).

OpenStudy (anonymous):

C=2r*pi=24pi \[\theta/360^{o}=(arclength)/C\]

OpenStudy (anonymous):

\[\theta/360^{o}=4\pi/24\pi=1/6\]\[\theta=360^{o}/6=60^{o}=\pi/3 radians\]